OpenStudy (anonymous):

If a horse does 4000 J of work over 20 m, how much force does the horse use?

OpenStudy (anonymous):

200 N

OpenStudy (anonymous):

F=W/d F=4000J/20m=200N

OpenStudy (anonymous):

Why do physics questions find their way in the math section ?? :D

OpenStudy (anonymous):

@anonymoustwo44 is absolutely right